OpenStudy (anonymous):

Find the exact value of sin(pi/12) = sin [(1/2)(pi/6)]

OpenStudy (jamesj):

http://openstudy.com/study#/updates/4f2588fde4b0a2a9c266f69f Now the third time you've asked. One way is to write \[ \sin(2x) = 2\sin x . \cos x = 2\sin x \sqrt{1-\sin^2 x} \] and set x = pi/12. You know what sin(pi/6) is. Then solve for sin x = sin(pi/12)